Calculate change in entropy of when 1 litre of water at 27℃is heated to 77℃

Grade:12th pass

1 Answers

Arun
4 years ago
 
 
w = 1 lit = 1 kg (approx)
hence mole n = 1000/18.5 = 54.054
 
now entropy change = n * cp ln (T2/ T1)
 
 = 54.054 * 4.2 * ln (360/300)
 
Hope it helps